OpenStudy (anonymous):

Mike and Kim invest $12,000 in equipment to print yearbooks for schools. Each yearbook costs $5 to print and sells for $15. How many yearbooks must they sell before their business breaks even? A. 800 yearbooks B. 1,200 yearbooks C. 2400 yearbooks D. 1,800 yearbooks

OpenStudy (anonymous):

i got D.

Parth (parthkohli):

Well, they get $10 profit. They need to get the $10 profits add up to 12000. So, 10 * 1200 = 12000 They need to sell 1200 yearbooks.

Parth (parthkohli):

It'd rather be option B.

Parth (parthkohli):

Or basically, 10x = 12000 x = 12000/10 = 1200

OpenStudy (anonymous):

@ParthKohli why 10 for the profit??

Parth (parthkohli):

Because they spend 5 dollars and then sell it for 15.

OpenStudy (anonymous):

I see tnx